find dy/dx y= (e)^x + (10)^x + (x)^x

find dy/dx
y= (e)^x + (10)^x + (x)^x

Grade:12

3 Answers

Rinkoo Gupta
9 years ago

Let y1=xx

Taking log on both sides, get

Logy1=xlogx

Diff w. r. to x ,get

1/y1. Dy1/dx=x.1/x+logx.1

Dy1/dx=y1[1+logx]

=xx[1+logx]

10^x = e^(x ln 10)

Applying the chain rule:dy/dx = e^(x ln 10) * ln 10 = 10^x ln 10

thus dy/dx=ex+10xln10 +xx(1+logx)

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty

Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty